OpenStudy (anonymous):

how do you convert 2 square miles to square meters? please show work

OpenStudy (bibby):

how many meters are in a mile?

OpenStudy (anonymous):

1609.34

OpenStudy (bibby):

Google says 1609.34m per mile Now we can use dimensional analysis like so.\[\large 1mi^2=\frac{1\cancel{mi}*\cancel{mi}}{}*\frac{1609.34m}{1\cancel{mi}}*\frac{1609.34m}{1\cancel{mi}}\]

OpenStudy (bibby):

I meant to have a 2 there. \[\large 2mi^2=\frac{2\cancel{mi}*\cancel{mi}}{}*\frac{1609.34m}{1\cancel{mi}}*\frac{1609.34m}{1\cancel{mi}}\]